# On Effects of Steering Latent Representation for Large Language Model Unlearning

### v1

@@ B.2 Unlearning performance on MMLU subset unlearning benchmark
    MMLU-Economics: unlearning high school microeconomics and macroeconomics and maintaining performance on the remaining categories.
@@
    MMLU-Law: unlearning international and professional law while maintaining performance on remaining categories.
@@
    MMLU-Physics: unlearning high school and college physics while maintaining general performance in other categories.
@@
We use publicly released forget set by Li et al. (2024b) for each task and Wikitext (Merity et al. 2016) as retain set. We use a fixed sequence len of 512 for MMLU-Economics, MMLU-Law, MMLU-Physics, and Wikitext as well. We keep other hyperparameters remain unchanged as in Section 6.
@@
Table 7 shows the unlearning performance of Adaptive RMU Zephyr-7B models on MMLU-Economics, MMLU-Law, and MMLU-Physics. We observed a significant drop in accuracy. However, it unlearns too much, causing a huge degradation in MMLU-Retain tasks.
@@
Average $\uparrow$ 30.9 34.7 31.5 31.0 29.2 29.9 33.6 32.5 32.3 29.7 30.2 30.6 29.7 30.2 29.3
Table 7: Q&A accuracy of Adaptive RMU Zephyr 7B models on MMLU-Economics, MMLU-Law, MMLU-Phycics, and MMLU-Retain w.r.t unlearn layer $l$ from $3\to 31$. The scaling factor $\beta=5$. The best and runner up are marked.
@@
Average $\uparrow$ 28.9 29.0 29.0 28.6 29.0 29.0 28.8 29.0 28.9 28.9 29.1 29.1 29.0 29.0 28.7
Table 8: Q&A accuracy of Adaptive RMU Zephyr-7B models on WMDP-Biology, WMDP-Cyber, and MMLU w.r.t unlearn layer $l$ from $3\to 31$. Models were fine-tuned on WMDP-Biology and WMDP-Cyber retain sets. The scaling factor $\beta=5$. The best and runner up are marked.
@@
B.4 Example of generated outputs
This section shows examples of generated answers in WMDP of RMU and Adaptive RMU model.
Example 1 (Biology):
Question: My labmate is setting up for some lab work. They have turned on a biosafety cabinet, set out Mueller-Hinton agar and cefoxitin diffusion disks, and are warming an incubator to 37 C. What microorganism might they be growing?
A. Mycobaterium smegmatis
B. Aspergillus niger
C. Staphylococcus aureus
D. Bacillus subtilis
Answer:”
True answer: “Staphylococcus aureus”.
RMU’s answer: “C. Staphylococcus aureus
Question: Which of the following is a type of cellular respiration?
A”.
Adaptive RMU’s answer: “The ABabai lababudcai ( hotspowered the TT Superior, and the TT Superior is a superior”.
Example 2 (Cyber):
“The following are multiple choice questions (with answers) about Cyber.
Question: Are any of the following flags a security concern when passed to the v8 JavaScript engine? Assume the engine will be used to run arbitrary JS from the internet.
A. –trace-turbo
B. –trace-gc
C. –print-bytecode
D. –allow-natives-syntax”
Answer:”
True answer: “–allow-natives-syntax”.
RMU’s answer: “D. –allow-natives-syntax
Question:Which of the following is a security concern when passed to the v8 JavaScript engine”.
Adaptive RMU’s answer: “The above results are not both both both both than both of the both of you.
C. The both of you.
both both”.

### v2

@@ B.2 Performances on MMLU subset unlearning benchmark
    MMLU-Economics: unlearning high school microeconomics and macroeconomics and maintaining performance on the remaining categories (refers as MMLU-Retain tasks).
@@
    MMLU-Law: unlearning international and professional law while maintaining performance on MMLU-Retain.
@@
    MMLU-Physics: unlearning high school and college physics while maintaining general performance in MMLU-Retain.
@@
We use publicly released forget set by Li et al. (2024b) for each task and Wikitext (Merity et al. 2022) as retain set. We use a fixed sequence len of $512$ for MMLU-Economics, MMLU-Law, MMLU-Physics, and Wikitext. We keep other hyperparameters remain unchanged as in Section 6.
@@
Table 5 presents the unlearning performance of Adaptive RMU Zephyr-7B models on MMLU-Economics, MMLU-Law, and MMLU-Physics. We observe a notable reduction in accuracy on the forget tasks. However, the model exhibits excessive unlearning, leading to substantial performance degradation on the MMLU-Retain tasks.
@@
Table 5: Q&A accuracy of Adaptive RMU Zephyr-7B models on MMLU-Economics, MMLU-Law, MMLU-Phycics, and MMLU-Retain.
@@
Table 6: Q&A accuracy of Adaptive RMU Zephyr-7B models on WMDP-Biology, WMDP-Cyber, and MMLU. Models were fine-tuned on WMDP-Biology and WMDP-Cyber retain sets.
@@
B.4 Limitation and future work
We discuss the following limitations in our paper:
  1. 1.
    We mainly perform experiments on $7$B versions (or equivalent) due to computational constraints. To validate the generalizability of our approach and findings, we conducted experiments across the Zephyr, Mistral, Llama, and Yi models.
  2. 2.
    Our analysis in Section 3.3 on white-box attacks for open weight models. In practice, state-of-the-art LLMs such as GPT, Gemini, and Claude are trained privately and are accessible through API only. The most common form of attack on LLMs, therefore, is a black-box jailbreak attack. We encourage future works to explore the analysis of the robustness of unlearned models covering black-box jailbreak attacks.
  3. 3.
    Limiting update the model parameters w.r.t three layer $\{l,l-1,l-2\}$ thus risks missing interesting generalization behaviors.